1)a If a sample of oxygen gas has a volume of 4.89 L at 20.0 °C and a pressure of 1.20 atm. If the temperature was increased to 56.0 °C then what would its resultant volume be (in L?) Assume the new pressure is 1.20 atm.

b At STP, 297.0 mL of a certain gas has a mass of 0.366 g. What is a possible identity of this gas?

Select one:

a. O3

b. Cl2

c. C2H6

d. O2

e. N2
Answer #1

volume         v1 = 4.89 l                 v2 = ?

temperature t1 = 20 c = 293k       t2 = 56c = 329k

pressure      p1 and p2 is 1.20 atm

no change in pressure so p1 = p2

then v1 / t1 = v2 / t2

         v2 = v1*t2 / t1

         v2 = 4.89 * 329 / 293

         v2 = 5.49 litres

answer new volume is = 5.5 litres

no of moles at STP as follows

at STP 1 mole occupies 22.4 litres

1mole----------------------> 22.4 litres

?   --------------------------> 0.297 litres

=>0.297/22.4 = 0.0132258 moles

answer is ( e ) = 0.366g / 0.0132258 moles   = 27.6 g / mol i.e = 28 g / mol i.e N2